### v1

# A note on thermodynamics

Abstract
The process of creating goods and services, measured by their value, is considered as a process of creating complexity. This makes it possible to consider the production system as an open thermodynamic system, and to develop a simple heuristic model for the production process. The model includes three production factors: the index of complexity of production equipment (physical capital $K$), human activity (labour $L$), and the substitutive capacity of equipment (substitutive work $P$). The latter is a contribution to economic theory from the thermodynamic approach, which also requires the introduction of technological characteristics of production equipment, such as labor requirement ($\overline{\lambda}$) and energy requirement ($\overline{\varepsilon}$), which indicate the amounts of labor and energy required to operate production equipment. By applying thermodynamic principles to the theory of production, we can understand how labour can be replaced by capital, and derive the production function in four equivalent but different formulations. Two of them are known and used by economists for interpretation the production phenomena; the thermodynamic approach gives some foundation for economic theory. The production function allows an unambiguously decompose of the growth rate of output according to the growth rates of production factors and technological level. The introduction of substitute work as a factor of production and technological features of capital expands planning and analyse of production processes.
Key words:
capital, labour, substitutive work, technological coefficients, production function
Introduction
Any production of consumables (food, clothes, buildings, transport, sewage systems, household appliances, machinery and so on) contains processes that convert natural resources into products for human usage. You can observe clay being turned into pots; or clay, sand and stone being transformed into buildings. Raw materials such as ore are turned into airplanes, which fly in the sky. This approach allows us to view the products we create as a result of transforming ’wild’ material into ’usable’ forms through work done by humans or machines (using energy from external sources such as wind, water, or coal). Any production system is being realized as a set of appliances for the matter transformations, and the system can be considered to be an open thermodynamic system. The simplest schematization of the production procces is shown in Fig. 1.
The aim of the production process is to obtain useful output $Y$, which is a monetory estimate of market value of created things. The production processes can be considered as processes of creation of value, and researchers have been engaged in finding the origins of this property, so called production factors. Historically, among the three factors listed in the input on Fig. 1, the aggregate human efforts — labour $L$ — was chosen by classical political economists [1, 2] as the sole source of value. That was a perfect solution until machines began to dominate workers’ efforts in production processes. To solve the problems that arise, the researchers incorporate another factor into their description. It was the amount of production equipment, also known as physical capital $K$, measured by its value, that was chosen as an aditional production factor.
Conventionally, the output of a production process, $Y$ is now understood as the result of activity of these two factors [3]. The first factor is labour efforts, $L$, which can be quantified, for example, in working hours per year. The second factor is physical capital, $K$, which represents the quantity of equipment used in production, so that the production of value
$Y=Y(L,K).$ (1)
The relation describes the effect of substitution; this means that the two production factors could substitute for each other. In other words, productive power is prescribed not only to labour, but also to the installed equipment; both of them appear to be the sources of value. The researchers are using the various specifications of function (1), adding some correcting assumptions, if necessary [3] .
Looking at Fig. 1, we are surprised; why was the amount of production equipment $K$, but not its work $P$, chosen as an argument for the production function. Indeed, how can labour be substituted by the amount of production equipment? One has to at least understand the mechanism of this substitution. Moreover, relation (1) could lead to some contradictions that are described in the literature [4].
In this paper, we propose considering the approach based on the concepts and principles of thermodynamics. The output of the production processes $Y$ is conventionally interpreted as market value estimated by consumers; however, this does not mean that it is a subjective quantity. The objective content of output appears to be complexity; the production of consumable goods can be interpreted as a process of creating useful complexity within the human environment; production activity moves boundaries between the artificial and natural environment. The created complexity $Y$ can be interpreted also as entropy that is diminishing in the human environment [5], which allows us to state that the natural measure of value is energy [5, 6]. Creation of complexity requires work that could be done by labor efforts, $L$, and/or machine work, $P$. Therefore, according to the thermodynamic laws, we can write
$Y=Y(L,P).$ (2)
Functions (1) and (2) represent two different approaches to describing production processes. There is a contradiction between these approaches, and a clear desire exists to find a unified description, which is the main goal of this paper. We believe that a thermodynamic approach can be useful in formulating the theory. Based on a simple diagram shown in Fig.1, we will develop in the next section a simple model that will allow us to determine the production function of a process in three forms. One form corresponds to relation (1), which describes the phenomenon of substitution, a well-known phenomena.
INPUTraw materials P machine work L labour efforts K production equipment $\overline{\lambda}$ labour req. $\overline{\varepsilon}$ energy req. OUTPUTheatpollutants Y useful output
Figure 1 Scheme of the production process
To produce a thing or service, apart from the production equipment $K$, one needs raw materials (ores, water, air, energy carriers and so on), worker efforts $L$ and machine work $P$. These factors are closely related to technology (the characteristics $\overline{\lambda}$ and $\overline{\varepsilon}$ ), embodied in the production equipment. The useful output of the production process $Y$ is a monetary estimate of the collection of useful things and services produced. Production processes are accompanied by heat and pollution emissions, but this is a separate issue; the ’useful’ side of production is addressed in the Letter.
Heuristic model of the production process
To determine specific forms of functions (1) and (2), we turn to the simple model of the production process, shown in Figure 1. The assembly of production equipment, that is a fixed embodiment of technology, commonly [3] characterised by physical capital $K$, measured in units of value [3] , can be considered as an open thermodynamic system. From a thermodynamic perspective, the variable $K$ is a characteristic of a non-equilibrium metha-stable state of the system, which appears to be the index of complexity. The production equipment is an essential part of modern enterprises, but it remains inactive until efforts by workers and energy input from external sources have been applied. The installed equipment determines how much labour $L$ and substitutive work $P$ will be required for the production process before it is activated. We assume that these factors of production are measured in energy units.
To fix the property of the capital to attract labour force and productive energy, we can , according to the previous speculation [7], introduce the main characteristics of the capital, which are labour and energy requirements,
$\lambda=\frac{\ d\,L}{\ d\,K},\quad\varepsilon=\frac{\ d\,P}{\ d\,K}.$ (3)
These quantities indicate the amounts of labour and productive energy required to operate a unit of production equipment.
It is convenient to deal with non-dimensional technological coefficients
$\overline{\lambda}(t)=\frac{K}{L}\,\frac{\ d\,L}{\ d\,K},\qquad\overline{\varepsilon}(t)=\frac{K}{P}\,\frac{\ d\,P}{\ d\,K}.$ (4)
If the quantity is less than one, it means that labor-saving or energy-saving technologies have been used at that moment of time. In a stationary case, the quantities do not change with time.
The quantities (4) characterize the basic technological level of production equipment. They determine the amount of required production factors. If the non-dimensional technological coefficients are considered to be given characteristics, then the relations (4) appear to be equations that determine the required amounts of production factors
$\frac{L}{L_{0}}=\left(\frac{K}{K_{0}}\right)^{\overline{\lambda}},\quad\frac{P}{P_{0}}=\left(\frac{K}{K_{0}}\right)^{\overline{\varepsilon}}.$ (5)
The symbols with zeros represent the values of the variables at a selected starting point.
The comparative efficiency of the use of production factors is characterized by a technological index, which is a combination of dimensionless technological coefficients
$\alpha=\frac{1-\overline{\lambda}}{\overline{\varepsilon}-\overline{\lambda}}.$ (6)
The technological coefficients and technological index are internal parameters of a production system. Variable $K$ represents the total amount (in monetary units) of production equipment, while the technological coefficients $\overline{\lambda}$ and $\overline{\varepsilon}$ indicate its quality. We assume that these three quantities are given as the global characteristics of production equipment.
Specification of the production function
In this section, we consider the output of a production process, as the market value of products created in a given point of time, $Y$, in relation to its value, $Y_{0}$ at an arbitrary starting point. This relative quantity can be expressed as a function of the production factors, $L$, $P$, and $K$, taken in relation to their values at the beginning of a period: $L_{0}$, $K_{0}$ and $P_{0}$. It will be convenient to use non-dimensional ratios of these variables further.
From a thermodynamic point of view, the true and only sources of value (as relatives of entropy) are labor expenses $L$ and the work of external energy sources $P$. Human efforts are applied directly, whereas substitutive work arises from the prior exploration of natural laws, the development of technological systems, and the manufacturing of production machinery, among other things. The interpretation of the production process allows us to view the output $Y$ as a function of two variables: the activity of workers — labour $L$ — and the work done by production equipment - substitutive work $P$, which depends, according to equation (5), on the amount of installed equipment $K$, so that we can write the following function
$Y=Y[L(K),P(K)].$ (7)
As the amounts of work, the production factors are interchangeable and, in this sense, they are equivalent in all respect. The Smith-Marx theory of labor value is supplemented by the law of substitution, which says that work of third-party forces of nature via production equipment replaces human efforts. Labor functions as labor plus work of equipment in complex in the production of value.
The general form of dependence of output on production factors can be determined by simple assumptions. Universality of production function requires that a proposed function can be applied not only to a specific case, but also to many different situations. For instance, the starting point can be arbitrary, but the shape of the function should not be influenced by this arbitrariness, which determines the unique power form of the function. Uniformity requires the arguments of the function to be special combinations of production factors. These requirements lead (details can be found in the author’s monograph [8, Chapter 6] to the expression for the production function in the form
$Y=Y_{0}\,\frac{L}{L_{0}}\left(\frac{L_{0}}{L}\frac{P}{P_{0}}\right)^{\alpha},\quad 0<\alpha<1,$ (8)
The index $\alpha$ — the technological index — in equation (8) determines the degree of substitution of labour with productive energy. It was shown in [8, Chapter 6; 9], considering the dynamics of the production factors, that the index $\alpha$ in equation (8) has the same form as it is defined by the formula (6); the level of technology development is included in the law of value production. Note that the labor $L$ and substitutive work of external energy sources $P$ is measured in energy units. Relation (8) was confirmed in many details in the author’s previous works [8, 9].
Relation (8) describes the effect of substitution labor efforts with machine work. Economists usually prefer to talk about substituting labor with capital [3]. Relations (5) allow us to represent the production of value (8) in terms of the variables $L$ and $K$. When we choose these variables, the resulting production function is as follows
$Y=Y_{0}\,\frac{L}{L_{0}}\left(\frac{L_{0}}{L}\left[\frac{K}{K_{0}}\right]^{\overline{\varepsilon}}\,\right)^{\alpha},\quad 0<\alpha<1.$ (9)
The recorded relation is a specific form of function (1), and represents the law of labor substitution by capital. An increase in output can be achieved through increasing relative labor expenses, or the relative amount of capital. The substitution coefficient depends on technological characteristics of production equipment. An analysis of substitution mechanisms shows that workers’ efforts to produce goods are substituted by the work of the production equipment, not by equipment itself.
Note that, due to the validity of equations (5), the functions (8) and (9) can be reduced to a simple relation
$Y=Y_{0}\,\frac{K}{K_{0}}.$ (10)
The proportionality of the output to the amount (in value) of capital is confirmed by observations [10, p. 4]. Existing deviations from this relationship may be due to different growth rates across industries [8, Chapter 9, section 9.3.2].
One of the relations (5) shows the amount of human effort for the presence of production equipment (physical capital $K$), so that the relations (9) and (10) are reduced to the form
$Y=Y_{0}\,\left(\frac{L}{L_{0}}\right)^{1/\overline{\lambda}}.$ (11)
This relation gives us the reason to say that labor remains, to use the words of Adam Smith, ”the only universal, as well as the only accurate measure of value, or the only standard by which we can compare the values of various goods at all times and in all places”. The law of the value production appears to be a nonlinear, generally speaking, function.
Expressions (8)-(11) represent four forms of production function as a consequence of a thermodynamic approach. The theory does not need to contain any arbitrary parameters to be consistent with empirical observations.
The marginal productivity of capital
As can be seen from the production process description, the useful output $Y$ is determined by three production factors, $L$, $P$ and $K$. Only two of them, $L$ and $P$, perform work; they are active factors from thermodynamic point of view. In contrast to this, capital $K$ appears to be the means, which allows the active factors to perform work, and should be considered as a passive factor of production.
Three forms of production function give rise to three different interpretations of processes of value production . Economists usually prefer to interpret the situations in terms of labor and capital [3]. Within the framework of this interpretation, the concept of the production power of capital has been formed, and the idea of capital productivity has become ingrained in the public consciousness. This applies not only to physical capital, but also to financial capital in various forms. For example, if we own shares in a company, we receive dividends; if we deposit money in a bank, we earn interest. Shares and money represent capital in its broadest sense.
Thermodynamic interpretation allows us to find an expression for capital productivity. To separate the individual contribution of production factors to the production of value we refer to the so-called marginal productivity of production factors. The active factors are characterized by the marginal productivities
$\beta=\frac{\partial Y}{\partial L},\quad\gamma=\frac{\partial Y}{\partial P}.$ (12)
Unlike factors $L$ and $P$, physical capital is not a source of value itself, but due to its role in production the ”marginal product of capital”,   can be formally calculated. After differentiating relation (7) we find that.
$\frac{d\,Y}{d\,K}=\beta\frac{\ d\,L}{\ d\,K}+\gamma\frac{\ d\,P}{\ d\,K}$ (13)
This expression shows that capital productivity is determined by marginal productivities of active production factors (12), and technological coefficients. One can see from equation (13) that the productive capacity of capital for the production of value is actually due to the productive capacity of labour and substitutive work.
The concluding remarks
In the paper, an interpretation of production process as process of creating complexity was proposed. In order to explain and describe increased labor productivity through mechanization, we need to take into account three fundamental factors of production: capital ($K$), labor ($L$), and substitute work ($P$). The latter is a contribution to economic theory from the thermodynamic approach, which also requires the introduction of technological characteristics of production equipment, such as labor requirement ($\overline{\lambda}$) and energy requirement ($\overline{\varepsilon}$), which indicate the amounts of labor and energy required to operate production equipment. By invoking thermodynamic principles in the theory of production processes, we can explain how labour is replaced by capital in production and derive a production function with four equivalent formulations (8) - (11). One can be satisfied that two forms (9) and (10) were known and were used by the economists for interpretation the production phenomena. In other words, the thermodynamic approach gives some foundation for relations of economic theory. The introduction of technological features of capital and substitute work as a factor of production expands the ability to plan and analyze the production. According to the theory, when describing the production of goods, qualitative technological attributes of capital matter more than its overall amount.
In conclusion, we should emphasize the fundamental importance of the law of value production. This law plays a crucial role in shaping the interpretation of production processes, and it serves as a foundation for determining how social goods are distributed. In essence, this law holds a fundamental position within the economic system and therefore deserves recognition as an essential economic principle. Finally, this question concerns the correct understanding of observable social phenomena.